OpenStudy (anonymous):

The first three terms of a sequence are given. 15,10,20/3,... Find the 6th term.

OpenStudy (across):

Let \(a_0=15\). Then\[a_{n+1}=\frac{2a_n}{3}\qquad n=0,1,\dots\]

OpenStudy (across):

Here, you're looking for \(a_5\).

OpenStudy (across):

That is,\[a_5=(2/3)(2/3)(2/3)(2/3)(2/3)a_0=\frac{32}{243}a_0=\frac{480}{243}\approx1.975308641975309.\]